# Particle Acceleration in Collisionless Magnetically Arrested Disks

### v1

@@ Abstract
We have highlighted the complex interplay between accretion-driven behavior, in the form of magnetic flux eruptions, and magnetospheric effects, with the included inverse-Compton scattering and pair production. The simulation demonstrate a considerable dynamic range in both mass density and magnetization. We identified the most prominent channels of particle acceleration, which are associated with pair production at the polar caps (pairs only) and to magnetic reconnection in the equatorial current sheet (pairs and ions). Additionally, we find that the jet is loaded with leptons solely created by pair production and virtually no ions are mixed into it. Significant mixing between disk and jet matter does occur at the disk-jet interface, primarily because of prominent KH-like modes that grow over time. These modes could lead to additional particle acceleration on the jet-disk boundary on larger scales [56]. The specifics of the gap remain debated [58, 59, 50, 60, 61, 33, 62], but from our simulation it becomes clear that the majority of highly energetic pairs are produced there, where two distinct channels can be identified for our modest multiplicities (Fig. 2). Furthermore, the BH spin is an influential factor for the gap characteristics that warrant more investigation. Future 3D studies will be even more informative on the kinetic effects related to the flux tubes themselves [e.g., 31].
@@
In this study, we presented the first numerical simulation that combines collisionless accretion (ion-electron disk) and magnetospheric (inverse-Compton scattering and pair creation) effects self-consistently using GRRPIC methods. It highlights the intricate dynamics of a quasi-steady MAD accretion state that is currently preferred for several of the most-observed supermassive BH systems [e.g., 7, 8]. The flux eruptions associated with these systems bring about strongly magnetized, largely vacated regions near the central BH, where strong particle acceleration processes are triggered — mostly in the equatorial current sheet and inner polar magnetospheric regions. The cyclic nature of the MAD flux eruptions makes them prime candidates for interpreting the periodic flares of high-energy emission observed from these systems [63, 64].
@@ Acknowledgements.
Supplemental Material for “Particle acceleration in Collisionless Magnetically Arrested Disks”
In this Supplemental Material, we provide more details on the initial conditions of the simulation as well as the procedure that underlies the spatially-decomposed spectrum calculations.
.1 Initial conditions and magnetic field initialization
Most of the details on the initial particle composition of the accretion flow have already been outlined in the main text, where the accretion environment contains three particle species (i.e., ion-pair plasma) that are organized in a ($8n_{\mathrm{GJ}}$, ion-electron plasma) wedge combined with a ($1n_{\mathrm{GJ}}$, electron-positron plasma) seeding sphere ($r<5r_{\mathrm{g}}$). For completeness, we explicitly show the initial conditions of $\beta=\sum_{s}\beta_{s}$ [65] and (cold) magnetization $\sigma_{\mathrm{c}}=B^{2}/\left[4\pi\sum_{s}m_{s}n_{s}\right]$ in Fig. 5. After an initial phase of strong pair-production (filling the empty space), we find that the accretion flow relaxes into a configuration dictated by the initialized magnetic field over $100$ $r_{\mathrm{g}}/c$.
@@
$A_{\phi}\propto re^{-r/400}\sin^{2}\theta,$ (1)
which is inspired by what is typically used to create a magnetically arrested disk (MAD) accretion state in GRMHD simulations [e.g., 31]. Customarily, however, the GRMHD magnetic field configuration is partly determined by the initialised hydrostatic disk solution [e.g., 66], which is technically not applicable for GRPIC methods so the magnetic field configuration inherently resembles a poloidal loop structure. The field was chosen so that $\bm{\Omega}_{\mathrm{h}}\cdot\bm{B}>0$. We ran the presented simulation for $1000$ $r_{\mathrm{g}}/c$ and eventually reached a quasi-steady accretion state with multiple flux eruption cycles.
@@
.2 Pair production mechanics and separation of scales
@@
In our simulation the $\gamma$-ray photons are evolved explicitly as a neutral fourth particle species, which comes about when a time-independent, uniform, and monochromatic soft background photon inverse Compton scatters of a relativistic lepton. The $\gamma$-rays are then able to pair-produce with the background radiation field. Specifics of the Monte Carlo procedure that lies at the basis of this can be found in Levinson and Cerutti [50], Crinquand et al. [33]. These dynamics are largely determined by the choice in fiducial optical depth $\tau_{0}=n_{0}\sigma_{\mathrm{T}}r_{\mathrm{g}}=50$, where $\sigma_{\mathrm{T}}$ is the Thompson cross-section and $n_{0}$ is the fiducial number density, which indicates that we operate in a relatively high optical depth regime that is likely significantly higher than is assumed for low-luminosity AGN (such as Sagittarius A and M87). The other fiducial parameters are the normalized magnetic field, $\tilde{B}_{0}=eB_{0}r_{\mathrm{g}}/m_{e}c^{2}=10^{4}$, and the normalized energy of the background photons, $\tilde{\epsilon}_{0}=\epsilon_{0}/m_{e}c^{2}=10^{-2}$, so that $\tilde{B}_{0}\tilde{\epsilon}_{0}=100$ at initialization (compared to $\tilde{B}_{0}\tilde{\epsilon}_{0}\sim 10^{6}$ for, e.g., Sagittarius A). This reflects an artificial shrinkage in the separation of scales, which is required to make the numerical study feasible [33]. Nevertheless, as becomes clear from Fig. 1 in the main text, the magnetic flux and corresponding magnetic field strength increases by a factor $20$, which pushes the simulation closer to a realistic separation of scales.
@@
We generally resolve the electron skin depth $d_{e}=\sqrt{m_{e}c^{2}/4\pi n_{\mathrm{GJ}}e^{2}}\sim r_{\mathrm{g}}/\sqrt{\tilde{B}_{0}}$ at a resolution of $N_{r}\times N_{\theta}=1440^{2}$ over the evaluated time, which implies that the ion skin depth $d_{i}=\sqrt{m_{i}/m_{e}}\,d_{e}$ is also well resolved. The reported simulation has a significant dynamical range in density $\rho$ and magnetic field strength $B$ (Fig. 2). During the quasi-steady accretion stages ($T\gtrsim 650\,r_{\mathrm{g}}/c$), particle phase-space is well sampled, with an O(60) number of electrons and positrons per cell. The total number of particles in the simulation is typically ${\sim}3\times 10^{8}$, where the majority consists of electrons and positrons.
.3 Spatially decomposed spectra calculations
@@
As discussed in the main text, we differentiate five distinct regions in which we evaluate the particle energy spectra; these regions are denoted as disk (dis), jet (jet), equatorial reconnection (rec) layer, jet sheath or jet-disk interface (she), and inner magnetosphere (inn). The decomposition of these regions is shown in Fig. 6. The mask is created with geometrical and dynamic criteria. For simplicity, we set the speed of light $c=1$ in the remainder of this section.
@@
Effectively, only the boundary between the disk and the jet sheath is determined by a dynamical criterion; when the quantity $\rho\langle\gamma\rangle>8\times 10^{5}$, we define it as disk, while when this criterion is not satisfied we assign it as jet sheath. Here, we define $\rho\langle\gamma\rangle=\sum_{s}m_{s}n_{s}\langle\gamma_{s}\rangle$, which contains the contribution of all species $s\in\{i,e^{-},e^{+}\}$. The remaining areas are the jet region (conical, [$\theta<\pi/4\vee\theta>3\pi/4$] & $r<5r_{\mathrm{g}}$, and then cylindrical, $r>5r_{\mathrm{g}}$ $\mathrm{X}<5\sin[\pi/4]$), the reconnection region ($\mathrm{X}<6r_{\mathrm{g}}$ $|\mathrm{Y}|<0.5r_{\mathrm{h}}$), and the inner magnetosphere ($r_{\mathrm{h}}<r<2r_{\mathrm{g}}\setminus\mathrm{\textit{rec}}$), which can be obtained from the figure itself. The geometrical areas are leading over the dynamically assigned areas. Lastly, we note that the spectra are only evaluated for $r\in[r_{\mathrm{h}},25r_{\mathrm{g}}]$.
@@
The spectra are created from reprocessed data of the original simulation that are outputted at a higher cadence and then put through the spatial-separation pipeline outlined in the previous paragraph. These individual spectra are then averaged to reduce the low-amplitude noise. Overall, the calculation of these spectra is a somewhat costly endeavor. We find that the construction of the spatial separation grid is robust as variations in the layout do not result in strongly differentiating spectra.
@@
Here, we comment on how $\gamma_{\mathrm{max}}$ is determined. Effectively, this is done with an empirical function defined according to $\max[(\gamma_{s}-1)^{p}\,f_{s}]$, where $f_{s}=dN/d\gamma_{s}$ is the particle energy spectrum and $p$ the empirical power law index that differs per particle species $s$. We found that a power law index $p\in[2,3.5]$ works well, where the highest $p$ was used for the electrons while the lowest was used for the positrons.
@@
Figure 6: Schematic outlining how the region-dependent spectra area are calculated. The starting, $\rho\langle\gamma\rangle$, quantity is time-averaged over a period of $T\in[726.85,730.33]\,r_{\mathrm{g}}/c$. This quantity is then blurred or smoothed to obtain a more smooth dissection between the disk (dis) and jet sheath (she) regions. Based on an empirically found criterion, one then makes the distinction. The remaining spatially decomposed regions are the equatorial reconnection (rec), jet (jet), and inner magnetosphere (inn) regions. The contours illustrate logarithmically scaled variations in the blurred $\rho\langle\gamma\rangle$ quantity.
@@
For completeness, we show the all components of the mapped spectra for the accretion states shown in the top and bottom panels of Fig. 2 in the main text. These panels correspond to the time instances $\mathrm{T}=730.97\,r_{\mathrm{g}}/c$ and $\mathrm{T}=881.23\,r_{\mathrm{g}}/c$, and are averaged over time windows $T\in[726.85,736.14]\,r_{\mathrm{g}}/c$ and T[874.45, $\mathrm{T}\in[874.45,\,887.50]\,r_{\mathrm{g}}/c$. For the state corresponding to $\mathrm{T}=881.23\,r_{\mathrm{g}}/c$, we find that $\gamma_{\mathrm{max},e^{+}}\gtrsim\gamma_{\mathrm{max},e^{-}}$, which coincides with a strong pair creation cascade that commences in the equatorial plane (forming a bubble structure, see also Fig. 8). Overall, we find that the creation of the associated gap occurs when the magnetic flux eruption is at its peak and the accretion flow has maximally retracted. The restoration of the accretion flow coincides with a kink mode (as discussed in the main text) where the strongest acceleration occurs. We note that the equatorial pair cascade (during which $\gamma_{\mathrm{max},e^{+}}\gtrsim\gamma_{\mathrm{max},e^{-}}$) is only a relatively rare occurrence and seems contingent on the strength and duration of the preceding flux eruption. We display the complete spectra of Fig. 4 of the main text in the left panels of Fig. 7 that corresponds to a general state where $\gamma_{\mathrm{max},i}<\gamma_{\mathrm{max},e^{+}}<\gamma_{\mathrm{max},e^{-}}$. During these period, the strongest acceleration occurs (for $e^{-}$) in the polar magnetospheric regions. The jet and jet sheath regions harbor less energetic electrons and positrons. For the ions, we see that moderate acceleration occurs in the sheath, which is connected to the regions containing KH-like modes. The strongest acceleration of ions is still associated with reconnection in the equatorial current sheet region.
@@
Occasionally, we find that the spectra, mostly for $e^{-}$, display a three-bumped structure, where the first corresponds to the base (thermal) disk component, the second originates in the reconnection region, and the third corresponds to the inner magnetospheric particle-creation dynamics, going from low to high Lorentz factor $\gamma$. Interestingly, the secondary reconnection-related bump is most intermittent; a late stage of the decay of this component is still visible in the right panels of Fig. 7. The intermittency of this reconnection component is to be expected, as the equatorial reconnection activity varies rapidly and is most directly affected by the MAD flux eruption dynamics.
.4 Inner light surface and gap dynamics
@@
with field orbital velocity $\Omega_{F}=-E_{\theta}/\sqrt{h}B^{r}$. Here, $h$ is the determinant of the spatial part of the metric and $g_{\phi\phi},g_{t\phi},g_{t\phi}$ are components of the Kerr metric [as outlined in detail in 67]. We note that the magnetic field configuration used here is new, but we find good correspondence to what was seen before with regard to the spark gap dynamics for purely magnetospheric studies [32, 33], except that some of the gap dynamics are largely dictated by the accretion component in our simulation. For instance, we see that the reinstating accretion flow also leaves an imprint on the inner light surface [33], where the otherwise smooth surface extrudes along the equatorial plane.
@@
In the next paragraphs, we further discuss the gap dynamics originally shown in Fig. 2 of the main text and display this region more explicitly in Fig. 8. As mentioned in the main text, much remains unclear about the formation, spatial extent, and location of the spark or electrostatic gap, which is characterized by significant quantities of unscreened electric fields denoted by $\bm{D}\cdot\bm{B}/B^{2}$ in Fig. 8, where $\bm{D}$ and $\bm{B}$ are the magnetic and electric fields as seen by fiducial observers [FIDOs; 67, 32]. In Fig. 8, we show the two distinct pair production channels found in our simulation, which are (i) located in the polar magnetospheric regions and (ii) in the equatorial plane (creating a bubble structure, as commented on in the main text). For completeness, we also show the inner light surface, which could be informative about the location of gap separation surface [between electron and positron; 33]. However, as discussed before [in, e.g., 50, 68, 62, 33], the gap is thought to be located close to the horizon at high optical depths [$\tau_{0}\geq 30$; 33]. Scenario (i) is consistent with the magnetospheric dynamics of a quasi-periodic, narrow gap opening near the polar axis. In panel 8c, we show that the unscreened electric field strength of this state which indicates that $|\bm{D}\cdot\bm{B}|/B^{2}\lesssim 10^{-1}$ near the gap, where the upper and lower hemisphere have an opposite polarity. In panel 8b, showing scenario (ii), we find that a bubble with highly energetic leptons forms, which is created because of the retracting accretion flow and a strong pair cascade that commences in the newly formed and large vacuum gap. As seen in panel 8d, the associated unscreened electric field is large with $|\bm{D}\cdot\bm{B}|/B^{2}\sim 1$, which explains the strong pair production cascade that commences in a relatively short time span. This type of equatorial pair cascade has not been seen before and is strongly related to the MAD flux eruption dynamics. It is also the moment when $\gamma_{\mathrm{max},e^{+}}\gtrsim\gamma_{\mathrm{max},e^{-}}$ and the highly-energetic positrons become more abundant. The strong kink mode corresponds to the restoring accretion flow and is also the separatrix between positive and negative unscreened electric fields.

### v2

@@ Abstract
We have highlighted the complex interplay between accretion-driven behavior, in the form of magnetic flux eruptions, and magnetospheric effects, with the included inverse-Compton scattering and pair production. We identified the most prominent channels of particle acceleration, which are associated with pair production at the polar caps (pairs only) and with magnetic reconnection in the equatorial current sheet (pairs and ions; Fig. 4). Additionally, we find that the jet is loaded with leptons solely created by pair production and virtually no ions are mixed into it (Fig. 3). Significant mixing between disk and jet matter occurs at the disk-jet interface, primarily because of prominent KH-like modes that grow over time. These modes could lead to additional particle acceleration on the jet-disk boundary on larger scales [56]. The specifics of the gap remain debated [58, 59, 50, 60, 61, 33, 62], but from our simulation it becomes clear that the majority of highly energetic pairs are produced there, where two distinct channels can be identified for our modest multiplicities (Fig. 2). Furthermore, the BH spin influences gap characteristics, warranting further investigation. Future 3D studies will be even more informative on the kinetic effects related to the flux tubes themselves [e.g., 31].
@@
In this study, we presented the first numerical simulation that combines collisionless accretion (ion-electron disk) and magnetospheric (inverse-Compton scattering and pair creation) effects self-consistently using GRRPIC methods. We highlight the intricate dynamics of a quasi-steady MAD accretion state that is currently preferred for several of the most-observed supermassive BH systems [e.g., 7, 8]. The flux eruptions associated with these systems bring about strongly magnetized, largely vacated regions near the central BH, where strong particle-acceleration processes are triggered — mostly in the equatorial current sheet and inner polar magnetospheric regions. The cyclic nature of MAD flux eruptions makes them prime candidates for interpreting the periodic flares of high-energy emission observed from these systems [63, 64].
@@ Acknowledgements.
Supplemental Material
In this Supplementary Material, we provide more details on the initial conditions of the simulation as well as the procedure that underlies the spatially decomposed spectrum calculations.
.1 Kerr-Schild coordinate system
Following generic numerical black hole modeling practice, we adopt horizon-penetrating Kerr-Schild coordinates (in the 3+1 decomposition [67] and closely following Komissarov [68]), which are defined according to
$ds^{2}=-\alpha^{2}dt^{2}+h_{ij}(dx^{i}+\beta^{i}dt)(dx^{j}+\beta^{j}dt),\hskip 20.00003pt\textrm{with}$ (1)
$h_{ij}=\begin{pmatrix}\Sigma\sin^{2}\theta/\rho^{2}&-a\sin^{2}\theta(1+z)&0\\ -a\sin^{2}\theta(1+z)&1+z&0\\ 0&0&\rho^{2}\end{pmatrix}.$
Here, $\alpha=1/\sqrt{1+z}$ is the lapse, $\beta=\left(0,z/(1+z),0\right)$ is the lapse, and $h_{ij}$ is the spatial metric with $\rho^{2}=r^{2}+a^{2}\cos^{2}\theta$, $\Delta=r^{2}-2Mr+a^{2}$, $\Sigma=(r^{2}+a^{2})^{2}-a^{2}\Delta\sin^{2}\theta$, and $z=2Mr/\rho^{2}$ for $i,j\in\{r,\theta,\phi\}$. For completeness, we note that $M$ is the black hole mass and that we have adopted geometrized units with $G=c=1$ here. For the non-geometrized version of the coordinates, one can simply replace $M$ with gravitational radius $r_{\mathrm{g}}=GM/c^{2}$. We note that the familiar form of the metric determinant is recovered as $\sqrt{-g}=\alpha\sqrt{h}$. All quantities are expressed in Kerr-Schild units throughout the main text and the supplemental material; figures are typically shown as a function of $\mathrm{X,\,Z}=r\cos\theta,\,r\sin\theta$. Lastly, we note that the underlying computational grid is uniformly spaced in $\hat{\theta}$ and logarithmically spaced in $\hat{r}$.
.2 Initial conditions, magnetic field initialization, and parameters exploration
Most of the details on the initial particle composition of the accretion flow have already been outlined in the main text, where the accretion environment contains three particle species (i.e., ion-pair plasma) that are organized in a ($8n_{\mathrm{GJ}}$, ion-electron plasma) wedge combined with a ($1n_{\mathrm{GJ}}$, electron-positron plasma) seeding sphere ($r<5r_{\mathrm{g}}$). We explicitly show the initial conditions of $\beta=\sum_{s}\beta_{s}$ [65] and (cold) magnetization $\sigma_{\mathrm{c}}=B^{2}/\left[4\pi\sum_{s}m_{s}n_{s}c^{2}\right]$ in Fig. 5. After an initial phase of strong pair-production (filling the empty space), we find that the accretion flow relaxes into a configuration dictated by the initialized magnetic field over $100$ $r_{\mathrm{g}}/c$.
@@
$A_{\phi}=B_{0}\left(\tfrac{r}{r_{\mathrm{g}}}\right)e^{-(r/r_{\mathrm{g}})/400}\sin^{2}\theta,$ (2)
which is inspired by what is typically used to create a magnetically arrested disk (MAD) accretion state in GRMHD simulations [e.g., 31]. Customarily, however, the GRMHD magnetic field configuration is partly determined by the initialized hydrostatic disk solution [e.g., 69], which is technically not applicable for GRPIC methods so the magnetic field configuration inherently resembles a poloidal loop structure. The field was chosen so that $\bm{\Omega}_{\mathrm{h}}\cdot\bm{B}>0$. We ran the presented simulation for $1000$ $r_{\mathrm{g}}/c$ and eventually reached a quasi-steady accretion state with multiple flux eruption cycles  [^2] For clarity, we further nuance what is meant by quasi-steady accretion state. MAD accretion states are characterized by magnetic flux eruption cycles that can strongly perturb the inflow of matter. This effect is reinforced in the relatively confining characteristics of axisymmetric simulations [e.g., 66]. So, when we speak of a quasi-steady state, we refer to the flattening over time of the $\dot{M}$ and $\Phi_{\mathrm{B}}$ timeseries shown in Fig. 1 of the main text..
For completeness, we note that while the contents of the main text and supplemental material are based on a single simulation, many exploratory runs have been undertaken. The most comparable simulations were performed with a lower ion mass ($m_{i}=50m_{e}$) and variations in initialized disk number densities ($8n_{\mathrm{GJ}}$ and $12n_{\mathrm{GJ}}$). The general findings for those simulations correspond well to what was reported for the simulation featured in the main text and supplemental material. These other simulations also show the formation of prominent Kelvin-Helmholtz (KH) eddies, as well as periodically repeating flux eruptions. Furthermore, most of the exploratory runs were conducted with a significantly smaller resolution ($N_{r}\times N_{\theta}=864^{2}$) for which we found similar behavior (albeit less well resolved).
@@
.3 Pair production mechanics and separation of scales
@@
In our simulation the $\gamma$-ray photons are evolved explicitly as a neutral fourth particle species, which comes about when a time-independent, uniform, and monochromatic soft background photon inverse Compton scatters of a relativistic lepton. The $\gamma$-rays are then able to pair-produce with the background radiation field. The details of the Monte Carlo procedure that lies at the basis of this can be found in Levinson and Cerutti [50] and Crinquand et al. [33]. These dynamics are largely determined by the choice in fiducial optical depth $\tau_{0}=n_{0}\sigma_{\mathrm{T}}r_{\mathrm{g}}=50$, where $\sigma_{\mathrm{T}}$ is the Thompson cross-section and $n_{0}$ is the fiducial number density, which indicates that we operate in a relatively high optical depth regime that is likely significantly higher than is assumed for low-luminosity AGN (such as Sagittarius A and M87). The other fiducial parameters are the normalized magnetic field, $\tilde{B}_{0}=eB_{0}r_{\mathrm{g}}/m_{e}c^{2}=10^{4}$, where the initial magnetic field strength is $B_{0}=10^{4}m_{e}c^{2}/|e|$ (see also the extended discussion in Parfrey et al. [32]). The normalized energy of the background photons, $\tilde{\epsilon}_{0}=\epsilon_{0}/m_{e}c^{2}=10^{-2}$, so that $\tilde{B}_{0}\tilde{\epsilon}_{0}=100$ at initialization (compared to $\tilde{B}_{0}\tilde{\epsilon}_{0}\sim 10^{6}$ for, e.g., Sagittarius A). This reflects an artificial shrinkage in the separation of scales, which is required to make the numerical study feasible [33]. Nevertheless, as becomes clear from Fig. 1 in the main text, the magnetic flux and corresponding magnetic field strength increase by a factor $20$ over the duration of the simulation, which indicates movement towards a more realistic separation of scales.
@@
We generally resolve the relativistic (and non-relativitic) electron skin depth $d_{e}=\sqrt{\langle\gamma_{e}\rangle m_{e}c^{2}/4\pi n_{\mathrm{GJ}}e^{2}}\sim r_{\mathrm{g}}/\sqrt{\tilde{B}_{0}}$ at a resolution of $N_{r}\times N_{\theta}=1440^{2}$ over the evaluated time, which implies that the ion skin depth $d_{i}=\sqrt{\langle\gamma_{i}\rangle m_{i}/\langle\gamma_{e}\rangle m_{e}}\,d_{e}$ is also well resolved. The reported simulation has a significant dynamical range in mass density and magnetic field strength (see Fig. 5 and Fig. 2 of the main text). During the quasi-steady accretion stages ($T\gtrsim 650\,r_{\mathrm{g}}/c$), particle phase space is well sampled, that is, the average number of electrons and positrons per cell is of O(60). The total number of particles in the simulation is typically ${\sim}3\times 10^{8}$, consisting primarily of electrons and positrons.
.4 Spatially decomposed spectra calculations
@@
As discussed in the main text, we differentiate five distinct regions in which we evaluate the particle energy spectra; these regions are denoted as jet (jet), jet sheath or jet-disk interface (she), equatorial reconnection (rec), inner (polar) magnetosphere (inn), and disk (dis) regions. The decomposition of these regions is shown in Fig. 6. The mask is created with geometrical and dynamic criteria. For simplicity, we set the speed of light $c=1$ in the remainder of this section.
@@
Effectively, only the boundary between the disk and the jet sheath is determined by a dynamical criterion; when the quantity $\zeta>1950$ (Fig. 6), we define it as disk, while when this criterion is not satisfied, we assign it as jet sheath. Here, we define $\rho\langle\gamma\rangle=\sum_{s}m_{s}n_{s}\langle\gamma_{s}\rangle$, which contains the contribution of all species $s\in\{i,e^{-},e^{+}\}$. The remaining areas are the jet region (conical, [$\theta<\pi/4\vee\theta>3\pi/4$] & $r<5r_{\mathrm{g}}$, and then cylindrical, $r>5r_{\mathrm{g}}$ $\mathrm{X}<5\sin[\pi/4]$), the reconnection region ($\mathrm{X}<6r_{\mathrm{g}}$ $|\mathrm{Y}|<0.5r_{\mathrm{h}}$), and the inner magnetosphere ($r_{\mathrm{h}}<r<2r_{\mathrm{g}}\setminus\mathrm{\textit{rec}}$), which can be obtained from the figure itself. The geometrical areas are leading over the dynamically assigned areas. We note that the spectra are only evaluated for $r\in[r_{\mathrm{h}},25r_{\mathrm{g}}]$.
@@
The spectra are created from reprocessed data of the original simulation that are outputted at a higher cadence and then put through the spatial-separation pipeline outlined in the previous paragraph. These individual spectra are then averaged to reduce the low-amplitude noise. The calculation of these spectra is a relatively costly endeavor. We verified that the resulting spectra remain consistent under (small) variations in the spatial decomposition mask.
@@
Here, we comment on how $\gamma_{\mathrm{max}}$ is determined. This is done by finding the Lorentz factor corresponding to the 80th percentile of the cumulative distribution function of the function $f_{s}=\gamma_{s}^{p}dN/d\gamma_{s}$, where $dN/d\gamma_{s}$ is the particle energy spectrum and $p$ the empirical power law index that differs per particle species $s$. For electrons and positrons, we use $p_{e^{\pm}}=1.5$. As the ions have a comparatively larger thermal component, we need to adopt a larger power law index to be most accurate, for which we choose $p_{i}=2.5$.
@@
Figure 6: Schematic showing how the spatially-dependent spectra are calculated. The quantity we evaluate is the normalized energy density $\zeta=\sum_{s}\langle\gamma_{s}\rangle\rho_{s}c^{2}/\rho_{\mathrm{GJ}}c^{2}$ with $\rho_{\mathrm{GJ}}=m_{e}n_{\mathrm{GJ}}$ and $\rho_{s}=m_{s}n_{s}$ for all particle species $s\in\{e^{-},e^{+},i\}$. This quantity is time-averaged over a period $T\in[726.85,730.33]\,r_{\mathrm{g}}/c$ and subsequently blurred to obtain a smooth dissection between the disk (dis) and jet sheath (she) regions. Based on the empirically found criterion (such as $\zeta>1950$), one then makes the distinction (for the disk region). The remaining spatially decomposed regions are the jet (jet), equatorial reconnection (rec), and inner (polar) magnetosphere (inn) regions. The contours illustrate logarithmically scaled variations in the blurred $\zeta$ quantity.
@@
For completeness, we show all the components of the mapped spectra for the accretion states shown in the top and bottom panels of Fig. 2 in the main text. These panels correspond to the time instances $\mathrm{T}=730.97\,r_{\mathrm{g}}/c$ and $\mathrm{T}=881.23\,r_{\mathrm{g}}/c$, and are averaged over time windows $T\in[726.85,736.14]\,r_{\mathrm{g}}/c$ and T[874.45, $\mathrm{T}\in[874.45,\,887.50]\,r_{\mathrm{g}}/c$. For the state corresponding to $\mathrm{T}=881.23\,r_{\mathrm{g}}/c$, we find that $\gamma_{\mathrm{max},e^{+}}\gtrsim\gamma_{\mathrm{max},e^{-}}$, which coincides with a strong pair creation cascade that commences in the equatorial plane (forming a bubble structure, seen clearly in the right panels of Fig. 8). Overall, we find that the creation of the associated gap occurs when the magnetic flux eruption is at its peak and the accretion flow has maximally retracted. The restoration of the accretion flow coincides with a kink mode (as discussed in the main text) where the strongest acceleration occurs. We note that the equatorial pair cascade (during which $\gamma_{\mathrm{max},e^{+}}\gtrsim\gamma_{\mathrm{max},e^{-}}$) is only a relatively rare occurrence and seems contingent on the strength and duration of the preceding flux eruption. We show the complete spectra of Fig. 4 of the main text in the left panels of Fig. 7 that correspond to a general state where $\gamma_{\mathrm{max},i}<\gamma_{\mathrm{max},e^{+}}<\gamma_{\mathrm{max},e^{-}}$. During these periods, the strongest acceleration occurs (for $e^{-}$) in the polar magnetospheric regions. The jet and jet sheath regions harbor less energetic electrons and positrons. For the ions, we see that moderate acceleration occurs in the sheath, which is connected to the regions containing KH-like modes. The strongest acceleration of ions is still associated with reconnection in the equatorial current sheet region.
@@
Occasionally, we find that the spectra, mostly for $e^{-}$, display a three-bumped structure, where the first corresponds to the (thermal) disk component, the second originates in the reconnection region, and the third corresponds to the inner magnetospheric particle-creation dynamics, going from low to high Lorentz factor $\gamma$, respectively. Interestingly, the secondary reconnection-related bump is most intermittent; a late stage of the decay of this component is still visible in the right panels of Fig. 7. The intermittency of this reconnection component is expected, as the equatorial reconnection activity varies rapidly and is most directly affected by the MAD flux eruption dynamics.
To assess the statistical significance of the spectra presented in the main text and this supplemental material, we report that the spectra (shown in Fig. 7) contain a total number of $n_{\mathrm{tot}}\approx 1.5\times 10^{8}$ particles. These are distributed among electrons, positrons, and ions with typical fractions of $f_{e^{-}}=0.5$, $f_{e^{+}}=0.4$, and $f_{i}=0.1$, respectively. The individual spectra are then combined and averaged over 160 time instances, evenly spread over the indicated time window, to mitigate the largest temporal variations. These values are most representative for the later stages of the simulation, of which we show the spectra. The ratios may seem skewed but we emphasize that not all particles have the same “numerical weight.” The disk-initialized electrons and ions represent a particle packet equivalent to 8 individual particles.
.5 Inner light surface and gap dynamics
@@
with field orbital velocity $\Omega_{F}=-E_{\theta}/\sqrt{h}B^{r}$. Here, $h$ is the determinant of the spatial part of the metric and $g_{\phi\phi},g_{t\phi},g_{t\phi}$ are components of the Kerr metric [as outlined in detail in 68]. We note that the magnetic field configuration used here is new, but we find good correspondence to what was seen before with regard to the spark gap dynamics for purely magnetospheric studies [32, 33], except that some of the gap dynamics are largely dictated by the accretion component in our simulation. For instance, we see that the reinstating accretion flow also leaves an imprint on the inner light surface [33], where the otherwise smooth surface extrudes along the equatorial plane.
@@
In the next paragraphs, we further discuss the gap dynamics originally shown in Fig. 2 of the main text and display this region more explicitly in Fig. 8. As mentioned in the main text, much remains unclear about the formation, spatial extent, and location of the spark or electrostatic gap, which is characterized by significant quantities of unscreened electric fields denoted by $\bm{D}\cdot\bm{B}/B^{2}$ in Fig. 8, where $\bm{D}$ and $\bm{B}$ are the electric and magnetic fields as seen by fiducial observers [FIDOs; 68, 32]. In Fig. 8, we show the two distinct pair production channels found in our simulation, which are (i) located in the polar magnetospheric regions and (ii) along the equatorial plane (creating a bubble structure, as commented on in the main text). For completeness, we also show the inner light surface, which could be informative about the location of gap separation surface [between electron and positron; 33]. However, as discussed before [in, e.g., 50, 71, 62, 33], the gap is thought to be located close to the horizon at high optical depth [$\tau_{0}\geq 30$; 33]. Scenario (i) is consistent with the magnetospheric dynamics of a quasi-periodic, narrow gap opening near the polar axis. In Fig. 8(c), we show the unscreened electric field strength of this state, which indicates that $|\bm{D}\cdot\bm{B}|/B^{2}\lesssim 10^{-1}$ near the gap, where the upper and lower hemisphere have an opposite polarity. In Fig. 8(b), showing scenario (ii), we find that a bubble-shaped equatorial region with highly energetic leptons forms, which is created because of the retracting accretion flow and a strong pair cascade that commences in the newly formed and large vacuum gap. As seen in Fig. 8(d), the associated unscreened electric field is large with $|\bm{D}\cdot\bm{B}|/B^{2}\sim 1$, which explains the strong pair production cascade that commences in a relatively short time span. This type of equatorial pair cascade has not been seen before and is strongly related to the MAD flux eruption dynamics. It also corresponds to the moment when $\gamma_{\mathrm{max},e^{+}}\gtrsim\gamma_{\mathrm{max},e^{-}}$ and highly-energetic positrons become more abundant. The equatorial relativistic drift-kink mode corresponds to the restoring accretion flow and is also the separatrix between positive and negative unscreened electric fields.